OpenStudy (anonymous):

Find the slope and the y-intercept of the line. Write the fractional answer in lowest terms y=-5/9x-4/5

OpenStudy (anonymous):

The slope of the line is the coefficient on the x variable (the number in front of the x) The y intercept is the value you get for y= when you plug in 0 for x.

OpenStudy (anonymous):

If you post what you get for those I will either confirm, or help figure out where you went wrong.

OpenStudy (anonymous):

so the slope is -4/5 and the y-intercept is -5/9

OpenStudy (anonymous):

Other way around.

OpenStudy (anonymous):

-5/9 is the number in front of the x variable. Therefore it is the slope. Plugging in 0 for x you get y = (-5/9)(0) - 4/5 = 0 - 4/5 = -4/5 So (x,y) = (0,-4/5) is the y intercept.